OpenStudy (anonymous):

Zach ran t times around the track at school, which is one-fourth mile. He ran a total of nine-tenths mile. Solve for t. four and four-ninths five-eighteenths three and three-fifths nine-fortieths

OpenStudy (btaylor):

How would you set up an equation to represent the situation? Think that every time he runs around the track, he travels an additional 1/4 mile.

OpenStudy (anonymous):

9/10 - 1/4 everytime?

OpenStudy (btaylor):

kinda. I would set it up as: \[\frac{1}{4} \times t = \frac{9}{10}\] Could you solve that?

OpenStudy (anonymous):

so i would divide 1/4 on both sides?

OpenStudy (btaylor):

yes, which is the same as multiplying by 4.

OpenStudy (anonymous):

it would be 3 3/5? thats what i got

OpenStudy (btaylor):

yep! good job!

OpenStudy (anonymous):

thank you :DDDD

OpenStudy (btaylor):

yw